OpenStudy (anonymous):

Anybody that can help with linear approximations and walk me through it step-by-step?

OpenStudy (anonymous):

ask a question

OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate sqrt64.3 as follows: Let f(x)= sqrtx The equation of the tangent line to f(x) at x=64 can be written in the form y=mx+b

OpenStudy (anonymous):

m=dy/dx which satisfy a point (64,8) you will get an eqn let say y=sqrtx dy/dx=1/2sqrtx at x=64 dy/dx=1/2sqrt64=1/16=m now when x=64 y=8 8=64/16+b b=4 y=x/16+4 now put x=64.3

OpenStudy (anonymous):

and get y

OpenStudy (anonymous):

What do I do after I get y?

OpenStudy (anonymous):

that value is approximated value of 64.3